### v2

@@ 3.4 Piloting
We first conducted one pilot study for each group with student assistants from our lab (PC1, PP1, PE1), who had not previously worked on privacy-related projects. PP1 had professional experience as a web developer. Based on their feedback, we made minor adjustments to the study tasks to make them more comprehensible. Further, as our pilot participants did not solve all the tasks within an initial estimation of three hours, we increased the time limit to five hours and pre-implemented non-privacy parts for Task~3.2 to further save programming time. Further, we switched the order of Tasks 3.2 and 3.2, since due to the pre-implemented parts in Task~3.2, it was shorter than Task~3.2. Lastly, based on participant feedback, we added Task~3.2 to help participants get started with the tool.
@@
We conducted a second round of piloting with all three groups (PC2, PP2, PE2). PC2 worked full-time as a software developer, PP2 worked part-time as a security consultant and part-time as a student assistant in our lab, and PE2, part-time as a senior software developer. PP2 was piloting the study as part of their employment in our lab. PC2 and PE2 were recruited from a pool of past study participants. Participants did not manage to solve all the tasks due to time issues and programming difficulties during Task~3.2, "Advertisements." To alleviate the time issues, we formulated clear finish-states in the tasks, avoiding participants spending time on the tasks for unnecessary functionalities (e.g., front-end design of the search functionality). For the programming difficulties, we added an existing email function for doctors and pointed the participants to this function to use as a function to adapt.
@@ 3.5 Participants
Recruiting professional software developers for a lab study was challenging due to lack of time, spread-out geographical locations, and high cost. We recruited participants through three different channels: (1) six developers from past computer science studies unrelated to privacy or data protection who agreed to receive future study invitations from our research group, (2) eight through advertising the study through “Kleinanzeigen“ [46], a German Craigslist [47] equivalent, and (3) 16 by snowball sampling.
We stopped recruitment when no willing participants registered for the study. All 94 interested participants were first presented with a short survey containing the consent form and an outline of the study. Participants were then screened for eligibility using the following criteria: 18 years of age or older, employed in the field of software development, experienced with JavaScript, and passed two out of three screening questions for programmers [48, 49]. To prevent participants from answering the questions through Google, we used the time limits for each question as recommended in [49].
Out of 94 participants who started the survey, 74 participants passed the screening criteria. All participants who met our requirements were directed to a calendar application to schedule a study appointment. Thirty of those decided to participate, with some of the remaining ones contacting the researchers to inform them they could not participate due to scheduling conflicts or travel distance. Participants were then randomly assigned to their group when they registered for the study.
We derived the compensation based on past research with software developers, where the average hourly compensation was $57$€ [50]. Applying this rate to our five-hour study resulted in a total compensation of $285$€. We added an additional $30$€ to cover travel expenses. All participants were compensated with $315$€ for their participation in the study.
The demographics of our participants can be found in Table II. Of the $30$ participants, $24$ identified as male and six as female. Participants were 30.8 years old on average (min: 18, md: 27 max: 56, $\sigma$: 10.17). Most (20) had a Bachelor’s (14) or Master’s degree (6). Five achieved a High School degree or equivalent, two had some college experience, two more had a vocational degree, and one had Grad School experience. They worked an average of $5.25$ years in their current field (min: 0.5, md: 3, max: 25, $\sigma$: 5.85). The majority were employed full-time (18), $7$ were self-employed, and $3$ worked part-time. Two participants preferred not to answer the question. The number of employees in the largest company participants had worked in ranged from $10$ to 260 000260000260\,000260 000.
@@ 3.6.1 Source Code
Functionality: Two researchers, R1 and R2, independently reviewed the source code the participants created during the study. They checked each task to see if it was functionally fulfilled. For Task~3.2, the “medical_conditions” table needed to be saved though the backup function. Task~3.2 was considered solved functionally when participants deleted data from the specific user. For Task~3.2, a doctor user needed to be able to enter a medical condition in a search field, which returned some data. Lastly, Task~3.2 was considered solved functionally if a mail could be sent through the system and appear in a MailCatcher.
Privacy: R1 and R2 went through the source code of the participants, noting if and how the participants considered data protection in their implementations. The findings were reviewed by R3. Task~3.2 did not require any considerations for privacy compliance. For Task~3.2, the remaining user data in the database was considered non-compliant. If the backup was changed, they checked if all data connected to the user was gone or anonymized. If it was, the solution was rated as compliant. If only some data was gone, they considered it a partially compliant solution. In Task~3.2, they first checked if personal information was given to the doctor, which might identify the patient (e.g., full names, addresses, emergency contacts). Second, they checked if a filtering method was implemented, ensuring doctors could only see data from their own patients or patients who had given permission to share their data. A solution was for a doctor not to be given access to personal information, but all patients’ health data was considered partially compliant. Lastly, for Task~3.2, they considered a solution compliant if a check for consent was implemented.
@@ 3.6.3 Interview
R2 conducted 11 interviews, R4 conducted 19. Transcriptions were done using Amberscript (GDPR-compliant), checked and approved by R2. Researchers R4 and R5 analyzed the interview transcripts using thematic analysis [51, 52]. First, the two researchers created a codebook based on two transcripts. The two researchers then coded the interviews individually, resolving conflicts through regular meetings and discussions (e.g., missing codes, concept clarification). For example, in these discussions, they added or merged codes introduced by each researcher in the final codebook. After both researchers had coded all the interviews, they resolved all conflicts, for example, differentiating purpose limitation and data minimization, through discussion. Final findings were reviewed with R3. We achieved 100% inter-coder agreement. The final codebook can be found in the supplementary material (see Section 7).
@@ 3.7 Positionality Statement
As suggested in prior work (e.g., [53, 54]) we provide researchers’ backgrounds involved in this work. Researchers specializing in Application Security (R1), Usable Security and Privacy (R2, R4, R5), and Data Protection (R3) conducted data collection and analysis. R3 was additionally a certified data protection officer with relevant work experience and served in expert communication for the expert group.
@@ 3.9 Limitations
Our study has limitations that need to be considered when interpreting the results. First, we present the results of a qualitative study. Participants were aware that they were participating in a study and may have behaved differently than in the real world. To address this issue, we used a role-playing technique to counter environmental bias. While we cannot claim generalizability, we provide a first step to understand a complex problem. Second, all participants were recruited in Germany and thus are not representative of all developers. Third, while participants were presented with programming screening questions, not all may have had the required experience level to solve all the tasks. Finally, a key limitation of our study is its focus on GDPR. While other legislations incorporate similar concepts (e.g., CCPA), the specifics of compliance may vary. For instance, CCPA allows companies to refuse a deletion request [2], whereas GDPR enforces stricter data erasure rights [33]. Similarly, health data sharing under HIPAA follows different minimization rules [55], and opt-out consent for email ads is valid under CCPA but not GDPR [34]. As a result, our findings may not be fully generalized to compliance tasks governed by other regulatory frameworks. Future research is needed to examine how these differences impact participant decision-making and task performance.
@@ 4 Results
Task~3.2 Task~3.2 Task~3.2 Task~3.2 Time
@@
The findings from the code assessment are presented in Table III. If the implementation was incomplete but participants at least tried to fulfill privacy requirements based on our criteria outlined in Section 3.2, we recognized participants’ awareness of privacy compliance as partially fulfilled. All participants fulfilled Task~3.2. For all four main tasks, participants could achieve a functional solution, resulting in 40 possible functional solutions for each group. As only three tasks had compliance issues to consider, each group could achieve a maximum of 30 compliant solutions. Special cases where participants did not follow the task as intended were excluded. For all tasks combined, while the participants in the different groups delivered a similar number of functional solutions (Control: 28/40, Prompt: 28/38, two special cases, Expert: 30/40), the number of privacy-compliant solutions differed. The Control group had the lowest number of compliant solutions (1/30 fully compliant, and 2/30 partially), suggesting that developers could be less likely to consider privacy if they are not explicitly prompted. Next was the prompted group (3/29 fully, 3/29 partly, one special case), with the expert group having the most compliant solutions (7/30 fully, 1/30 partly). While only 4 participants in the expert group contacted the experts, the support of the experts could have potentially supported the participants enough to improve compared to the prompted group.
4.1 Task~3.2
4.1.1 Source Code Analysis for Task~3.2
Almost all participants successfully fulfilled Task~3.2. They were capable of identifying and addressing the backup issue, indicating competency in basic application maintenance tasks. Only one participant (P4) failed to complete the task, experiencing potential challenges or misunderstandings in the backup process. Another participant (E10) partially fulfilled the task by identifying the line of code responsible for excluding the "medical_condition" table from the backup. However, instead of fully addressing the issue, the participant opted to comment out the exclusion code, which led to errors during the backup process. Only one prompted participant did not provide a solution, and another expert participant partially fulfilled it for Task~3.2. {summary}Task 1 Overall, the task posted little challenges for the participants.
4.1.2 In Their Words: Interview Insights for Task~3.2
Six participants explicitly mentioned they saw no relevant issues with privacy regarding Task~3.2. Four others saw issues with the developers being able to access the data and suggested access control measures or permission systems. Three suggested some form of database obfuscation, such that the data would not be easily readable by developers:
%That the names should be not plain readable, the email addresses not. And the other, like more the medicines and so on.m — P7
@@
%Maybe like, yes, not backing up the whole table, but only extract parts from them that are needed for the functionality of the service, but not on the personal data, or ask for permission.m — E3
4.2 Task~3.2
4.2.1 Source Code Analysis for Task~3.2
Twenty-five participants functionally fulfilled Task~3.2 by successfully deleting all user data from the database. However, only four considered the backups, such as participants E6 and E9, who ensured that user data was also removed from backups, effectively mitigating privacy risks associated with retaining user data. Twenty-seven participants failed to fully address the privacy aspect of Task~3.2. While they deleted user data from the database, they missed the backups, posing ongoing privacy risks. Further, one participant deleted all existing backups to comply with the data deletion request, inadvertently removing important data backup measures. While this removes the data of the patient who requested deletion, it is obviously undesirable, as it reverses the benefits of a backup for all other data, potentially leading to data loss.
@@
Notably, the two participants in the expert group (E6 and E9) successfully fulfilled the task in terms of privacy by ensuring that user data was removed not only from the database but also from backups. Both participants made use of the chat functionality to contact the expert. However, only E6 asked if data needed to be deleted in the backups, with E9 only notifying the expert that they had done so. {summary}Task 2 While most participants were able to solve the task, privacy was rarely considered, and sensitive user data often remained in the application.
4.2.2 In Their Words: Interview Insights for Task~3.2
Concerning the deletion request in Task~3.2, nineteen participants were aware that data with dependencies must also be deleted. Ten of the participants explicitly mentioned that data in the backup would be relevant and had to be deleted in addition to data in the live database. However, only two did so. Fourteen participants mentioned deleting the data manually. Twelve also suggested an automated deletion process for such cases. Further, six participants mentioned that they, as developers, should not be able to see or edit the application users’ data and argued that some form of access control should be in place or data should be obfuscated in some form. Two participants, however, recognized the task as a deletion request, a right under the GDPR. Four participants also mentioned specific considerations on the deletion, such as timely deletion or keeping proof of deletion. However, two participants believed the user data was important and should not be deleted, which could be a violation if the data is not sufficiently anonymized:
%The second task was to delete user data. And from my perspective, this user data shouldn’t be deleted because it could be important information for other parties. Not the user itself but for other parties. And yeah, I wrote down that it’s not a good idea to delete important user data. It would be better to flag the user as deleted and make it possible for some information that has to be stored to be accessible in the future. But I deleted the user anyway.m — C1
@@
4.3 Task~3.2
4.3.1 Source Code Analysis for Task~3.2
@@
Twenty participants successfully fulfilled the functional aspects of Task~3.2 but failed to adequately address privacy concerns, resulting in unfiltered searches across the entire database table. This oversight means that doctors could access confidential patient information without proper authorization.
@@
Six participants fully integrated privacy considerations into their codebase, while three more partially addressed privacy concerns. Participants P2 and P3 partially fulfilled the privacy aspect of Task~3.2 by refraining from displaying any patient data during searches. However, their implementation allowed doctors to search across all entries in the "medical_condition" table without considering whether the medical conditions belong solely to their own patients. Although patient identities were safeguarded, doctors could still access comprehensive information about medications and medical conditions without knowing which patient they belonged to. Even if patient identities remain undisclosed, they allow access to potentially identifying data. Five participants (C2, P5, P6, E3, E4) demonstrated a proactive approach by limiting search options based on patient consent, thereby limiting doctors’ unauthorized access to medical records. This approach is consistent with existing application functionality where patients grant permission to access data. They displayed the patient’s first and last name and the needed information.
@@
P6 took a privacy-conscious approach by not only limiting search results based on patient consent but also anonymizing the data to increase confidentiality. By excluding patient names from the results and only displaying medical conditions and allergies, the participant prioritized maintaining privacy without compromising functionality. Similarly, participant E1 limited the search to patients who had previously scheduled an appointment with the respective doctor. This approach improves data protection by ensuring that data access is only granted to patients with previous interactions, thereby minimizing the risk of unauthorized information retrieval. Six control, seven prompted, and seven expert participants delivered functional solutions. In the control group, one participant provided a compliant solution, compared to two partially compliant and two fully compliant solutions in the prompted group. The participants in the expert group delivered three compliant and two partially compliant solutions. {summary}Task 3 While functionality was not an issue in this task, most participants did not attempt a privacy-compliant solution or failed to submit a privacy-compliant implementation.
4.3.2 In Their Words: Interview Insights for Task~3.2
@@
For Task~3.2, the implementation of a search function for doctors, twenty-eight participants came to the conclusion that patient data should not simply be available for all doctors. Fifteen participants mentioned some form of access control measures for doctors:
@@
%Yeah, I don’t think that I should query all the users and be able to see all the medical information from all users. All like, even if I’m the doctor of someone else, I can see the like, I can see a person As if I’m a person Bs doctor, I can see person As private information. That shouldn’t be possible, I think.m — C5
@@
%The deal here is to treat a certain medical condition and additional data like the first name, last name, birthday, I don’t know, blood type is not required in this case. I basically would only give back treatments or medications that have worked positively for this kind of condition in the past. And the rest is not valuable or not relevant for this situation, I would say.m — P5
@@
%I mean, I don’t know, the whole function, like the whole function seems to be a bit, not really privacy-protecting.m — C8
@@
%Technically, it’s possible. And I’m not sure if it’s enough, but at least one has to filter whether for each single user, one has the permissions to read it given by the user, and I would think that the user especially has to allow this kind of use.m — C2
4.4 Task~3.2
4.4.1 Source Code Analysis for Task~3.2
Thirteen participants successfully realized the functionality of Task~3.2. Out of those, however, none did so in a privacy-compliant way. Several noteworthy implementation aspects that did not quite fit the scoring scheme will be described below. They are denoted as “Special Case” (SC) in Table III. One participant (P3) misunderstood Task~3.2 and added the email functionality to the doctor’s account. Similarly, P8 did implement this task in a way that did not fit the task description either. Instead of sending it to a list of people, the admin has to type each receiver’s email by hand. Neither E1 nor E9 did finish the task from a functional point of view. However, the aspects implemented by either highlight that both wanted to check user consent prior to sending e-mails. E9 added a flag to the user’s profile page to opt into e-mail notifications, whereas E1 added a filter for user consent prior to sending e-mails.
Overall, for Task 4, five control participants delivered a functioning solution, and three prompted participants implemented a functioning solution and two special cases. Five expert participants delivered a functioning solution for the last task. No participant in the control group or the prompted group delivered a privacy-compliant solution for the fourth task. In contrast, two participants in the expert group were able to implement a compliant solution. {summary}Task 4 While the number of functioning solutions was similar for the groups, only the expert group was able to fulfill the privacy requirements.
4.4.2 In Their Words: Interview Insights for Task~3.2
During Task~3.2, twenty-two participants mentioned consent as the main concept to create a compliant solution:
%Yes, the parties have all consented to their email addresses being stored there. The question is whether they have also agreed to allow advertising to be sent out.m — E8
@@
%That shouldn’t be, that shouldn’t be allowed, I think. Especially that I used email addresses that were for doctor purposes to have for advertisement purposes.m — C5
Similar to Task~3.2, five had concerns with the task itself, seeing it as violating the users’ privacy, and were unsure if a compliant solution was possible:
%They’ve probably never even, have never even agreed to receive advertisement emails, so I think the entire task is probably faulty.m — P6
@@
%They are allowed to send the email data. So in this case, I think, this should be okay to send the patients emails.m — P9

### v3

@@ 3.4 Piloting
We first conducted one pilot study for each group with student assistants from our lab (PC1, PP1, PE1), who had not previously worked on privacy-related projects. PP1 had professional experience as a web developer. Based on their feedback, we made minor adjustments to the study tasks to make them more comprehensible. Further, as our pilot participants did not solve all the tasks within an initial estimation of three hours, we increased the time limit to five hours and pre-implemented non-privacy parts for Task 3.2 to further save programming time. Further, we switched the order of Tasks 3.2 and 3.2, since due to the pre-implemented parts in Task 3.2, it was shorter than Task 3.2. Lastly, based on participant feedback, we added Task 3.2 to help participants get started with the tool.
@@
We conducted a second round of piloting with all three groups (PC2, PP2, PE2). PC2 worked full-time as a software developer, PP2 worked part-time as a security consultant and part-time as a student assistant in our lab, and PE2, part-time as a senior software developer. PP2 was piloting the study as part of their employment in our lab. PC2 and PE2 were recruited from a pool of past study participants. Participants did not manage to solve all the tasks due to time issues and programming difficulties during Task 3.2, "Advertisements." To alleviate the time issues, we formulated clear finish-states in the tasks, avoiding participants spending time on the tasks for unnecessary functionalities (e.g., front-end design of the search functionality). For the programming difficulties, we added an existing email function for doctors and pointed the participants to this function to use as a function to adapt.
@@ 3.5 Participants
Recruiting professional software developers for a lab study was challenging due to lack of time, spread-out geographical locations, and high cost. We recruited participants through three different channels: (1) six developers from past computer science studies unrelated to privacy or data protection who agreed to receive future study invitations from our research group, (2) eight through advertising the study through “Kleinanzeigen“ [kleinanzeigen], a German Craigslist [craigslist] equivalent, and (3) 16 by snowball sampling.
We stopped recruitment when no willing participants registered for the study. All 94 interested participants were first presented with a short survey containing the consent form and an outline of the study. Participants were then screened for eligibility using the following criteria: 18 years of age or older, employed in the field of software development, experienced with JavaScript, and passed two out of three screening questions for programmers [danilova2021you, danilova2022testing]. To prevent participants from answering the questions through Google, we used the time limits for each question as recommended in [danilova2022testing].
Out of 94 participants who started the survey, 74 participants passed the screening criteria. All participants who met our requirements were directed to a calendar application to schedule a study appointment. Thirty of those decided to participate, with some of the remaining ones contacting the researchers to inform them they could not participate due to scheduling conflicts or travel distance. Participants were then randomly assigned to their group when they registered for the study. We derived the compensation based on past research with software developers, where the average hourly compensation was 5757€ [naiakshina2020on]. Applying this rate to our five-hour study resulted in a total compensation of 285285€. We added an additional 3030€ to cover travel expenses. All participants were compensated with 315315€ for their participation in the study.
The demographics of our participants can be found in Table II. Of the $30$ participants, $24$ identified as male and six as female. Participants were 30.8 years old on average (min: 18, md: 27 max: 56, $\sigma$: 10.17). Most (20) had a Bachelor’s (14) or Master’s degree (6). Five achieved a High School degree or equivalent, two had some college experience, two more had a vocational degree, and one had Grad School experience. They worked an average of $5.25$ years in their current field (min: 0.5, md: 3, max: 25, $\sigma$: 5.85). The majority were employed full-time (18), $7$ were self-employed, and $3$ worked part-time. Two participants preferred not to answer the question. The number of employees in the largest company participants had worked in ranged from $10$ to 260 000260\,000.
@@ 3.6.1 Source Code
Functionality: Two researchers, R1 and R2, independently reviewed the source code the participants created during the study. They checked each task to see if it was functionally fulfilled. For Task 3.2, the “medical_conditions” table needed to be saved though the backup function. Task 3.2 was considered solved functionally when participants deleted data from the specific user. For Task 3.2, a doctor user needed to be able to enter a medical condition in a search field, which returned some data. Lastly, Task 3.2 was considered solved functionally if a mail could be sent through the system and appear in a MailCatcher.
Privacy: R1 and R2 went through the source code of the participants, noting if and how the participants considered data protection in their implementations. The findings were reviewed by R3. Task 3.2 did not require any considerations for privacy compliance. For Task 3.2, the remaining user data in the database was considered non-compliant. If the backup was changed, they checked if all data connected to the user was gone or anonymized. If it was, the solution was rated as compliant. If only some data was gone, they considered it a partially compliant solution. In Task 3.2, they first checked if personal information was given to the doctor, which might identify the patient (e.g., full names, addresses, emergency contacts). Second, they checked if a filtering method was implemented, ensuring doctors could only see data from their own patients or patients who had given permission to share their data. A solution was for a doctor not to be given access to personal information, but all patients’ health data was considered partially compliant. Lastly, for Task 3.2, they considered a solution compliant if a check for consent was implemented.
@@ 3.6.3 Interview
R2 conducted 11 interviews, R4 conducted 19. Transcriptions were done using Amberscript (GDPR-compliant), checked and approved by R2. Researchers R4 and R5 analyzed the interview transcripts using thematic analysis [braun2006using, clarke2015thematic]. First, the two researchers created a codebook based on two transcripts. The two researchers then coded the interviews individually, resolving conflicts through regular meetings and discussions (e.g., missing codes, concept clarification). For example, in these discussions, they added or merged codes introduced by each researcher in the final codebook. After both researchers had coded all the interviews, they resolved all conflicts, for example, differentiating purpose limitation and data minimization, through discussion. Final findings were reviewed with R3. We achieved 100% inter-coder agreement. The final codebook can be found in the supplementary material (see Section 7).
@@ 3.7 Positionality Statement
As suggested in prior work (e.g., [Sannon2022PrivacyResearch, Ortloff2023DifferentResearchers]) we provide researchers’ backgrounds involved in this work. Researchers specializing in Application Security (R1), Usable Security and Privacy (R2, R4, R5), and Data Protection (R3) conducted data collection and analysis. R3 was additionally a certified data protection officer with relevant work experience and served in expert communication for the expert group.
@@ 3.9 Limitations
Our study has limitations that need to be considered when interpreting the results. First, we present the results of a qualitative study. Participants were aware that they were participating in a study and may have behaved differently than in the real world. To address this issue, we used a role-playing technique to counter environmental bias. While we cannot claim generalizability, we provide a first step to understand a complex problem. Second, all participants were recruited in Germany and thus are not representative of all developers. Third, while participants were presented with programming screening questions, not all may have had the required experience level to solve all the tasks. Finally, a key limitation of our study is its focus on GDPR. While other legislations incorporate similar concepts (e.g., CCPA), the specifics of compliance may vary. For instance, CCPA allows companies to refuse a deletion request [CCPA], whereas GDPR enforces stricter data erasure rights [gdprart17]. Similarly, health data sharing under HIPAA follows different minimization rules [HIPAAmin], and opt-out consent for email ads is valid under CCPA but not GDPR [gdprccpacomp]. As a result, our findings may not be fully generalized to compliance tasks governed by other regulatory frameworks. Future research is needed to examine how these differences impact participant decision-making and task performance.
@@ 4 Results
Task 3.2 Task 3.2 Task 3.2 Task 3.2 Time
@@
The findings from the code assessment are presented in Table III. If the implementation was incomplete but participants at least tried to fulfill privacy requirements based on our criteria outlined in Section 3.2, we recognized participants’ awareness of privacy compliance as partially fulfilled. All participants fulfilled Task 3.2. For all four main tasks, participants could achieve a functional solution, resulting in 40 possible functional solutions for each group. As only three tasks had compliance issues to consider, each group could achieve a maximum of 30 compliant solutions. Special cases where participants did not follow the task as intended were excluded. For all tasks combined, while the participants in the different groups delivered a similar number of functional solutions (Control: 28/40, Prompt: 28/38, two special cases, Expert: 30/40), the number of privacy-compliant solutions differed. The Control group had the lowest number of compliant solutions (1/30 fully compliant, and 2/30 partially), suggesting that developers could be less likely to consider privacy if they are not explicitly prompted. Next was the prompted group (3/29 fully, 3/29 partly, one special case), with the expert group having the most compliant solutions (7/30 fully, 1/30 partly). While only 4 participants in the expert group contacted the experts, the support of the experts could have potentially supported the participants enough to improve compared to the prompted group.
4.1 Task 3.2
4.1.1 Source Code Analysis for Task 3.2
Almost all participants successfully fulfilled Task 3.2. They were capable of identifying and addressing the backup issue, indicating competency in basic application maintenance tasks. Only one participant (P4) failed to complete the task, experiencing potential challenges or misunderstandings in the backup process. Another participant (E10) partially fulfilled the task by identifying the line of code responsible for excluding the "medical_condition" table from the backup. However, instead of fully addressing the issue, the participant opted to comment out the exclusion code, which led to errors during the backup process. Only one prompted participant did not provide a solution, and another expert participant partially fulfilled it for Task 3.2. Overall, the task posted little challenges for the participants.
4.1.2 In Their Words: Interview Insights for Task 3.2
Six participants explicitly mentioned they saw no relevant issues with privacy regarding Task 3.2. Four others saw issues with the developers being able to access the data and suggested access control measures or permission systems. Three suggested some form of database obfuscation, such that the data would not be easily readable by developers:
That the names should be not plain readable, the email addresses not. And the other, like more the medicines and so on. — P7
@@
Maybe like, yes, not backing up the whole table, but only extract parts from them that are needed for the functionality of the service, but not on the personal data, or ask for permission. — E3
4.2 Task 3.2
4.2.1 Source Code Analysis for Task 3.2
Twenty-five participants functionally fulfilled Task 3.2 by successfully deleting all user data from the database. However, only four considered the backups, such as participants E6 and E9, who ensured that user data was also removed from backups, effectively mitigating privacy risks associated with retaining user data. Twenty-seven participants failed to fully address the privacy aspect of Task 3.2. While they deleted user data from the database, they missed the backups, posing ongoing privacy risks. Further, one participant deleted all existing backups to comply with the data deletion request, inadvertently removing important data backup measures. While this removes the data of the patient who requested deletion, it is obviously undesirable, as it reverses the benefits of a backup for all other data, potentially leading to data loss.
@@
Notably, the two participants in the expert group (E6 and E9) successfully fulfilled the task in terms of privacy by ensuring that user data was removed not only from the database but also from backups. Both participants made use of the chat functionality to contact the expert. However, only E6 asked if data needed to be deleted in the backups, with E9 only notifying the expert that they had done so. While most participants were able to solve the task, privacy was rarely considered, and sensitive user data often remained in the application.
4.2.2 In Their Words: Interview Insights for Task 3.2
Concerning the deletion request in Task 3.2, nineteen participants were aware that data with dependencies must also be deleted. Ten of the participants explicitly mentioned that data in the backup would be relevant and had to be deleted in addition to data in the live database. However, only two did so. Fourteen participants mentioned deleting the data manually. Twelve also suggested an automated deletion process for such cases. Further, six participants mentioned that they, as developers, should not be able to see or edit the application users’ data and argued that some form of access control should be in place or data should be obfuscated in some form. Two participants, however, recognized the task as a deletion request, a right under the GDPR. Four participants also mentioned specific considerations on the deletion, such as timely deletion or keeping proof of deletion. However, two participants believed the user data was important and should not be deleted, which could be a violation if the data is not sufficiently anonymized:
The second task was to delete user data. And from my perspective, this user data shouldn’t be deleted because it could be important information for other parties. Not the user itself but for other parties. And yeah, I wrote down that it’s not a good idea to delete important user data. It would be better to flag the user as deleted and make it possible for some information that has to be stored to be accessible in the future. But I deleted the user anyway. — C1
@@
4.3 Task 3.2
4.3.1 Source Code Analysis for Task 3.2
@@
Twenty participants successfully fulfilled the functional aspects of Task 3.2, but fifteen of them failed to adequately address privacy concerns, resulting in unfiltered searches across the entire database table. This oversight means that doctors could access confidential patient information without proper authorization.
@@
Six participants fully integrated privacy considerations into their codebase, while three more partially addressed privacy concerns. Participants P2 and P3 partially fulfilled the privacy aspect of Task 3.2 by refraining from displaying any patient data during searches. However, their implementation allowed doctors to search across all entries in the "medical_condition" table without considering whether the medical conditions belong solely to their own patients. Although patient identities were safeguarded, doctors could still access comprehensive information about medications and medical conditions without knowing which patient they belonged to. Even if patient identities remain undisclosed, they allow access to potentially identifying data. Five participants (C2, P5, P6, E3, E4) demonstrated a proactive approach by limiting search options based on patient consent, thereby limiting doctors’ unauthorized access to medical records. This approach is consistent with existing application functionality where patients grant permission to access data. They displayed the patient’s first and last name and the needed information.
@@
P6 took a privacy-conscious approach by not only limiting search results based on patient consent but also anonymizing the data to increase confidentiality. By excluding patient names from the results and only displaying medical conditions and allergies, the participant prioritized maintaining privacy without compromising functionality. Similarly, participant E1 limited the search to patients who had previously scheduled an appointment with the respective doctor. This approach improves data protection by ensuring that data access is only granted to patients with previous interactions, thereby minimizing the risk of unauthorized information retrieval. Six control, seven prompted, and seven expert participants delivered functional solutions. In the control group, one participant provided a compliant solution, compared to two partially compliant and two fully compliant solutions in the prompted group. The participants in the expert group delivered three compliant and two partially compliant solutions. While functionality was not an issue in this task, most participants did not attempt a privacy-compliant solution or failed to submit a privacy-compliant implementation.
4.3.2 In Their Words: Interview Insights for Task 3.2
@@
For Task 3.2, the implementation of a search function for doctors, twenty-eight participants came to the conclusion that patient data should not simply be available for all doctors. Fifteen participants mentioned some form of access control measures for doctors:
@@
Yeah, I don’t think that I should query all the users and be able to see all the medical information from all users. All like, even if I’m the doctor of someone else, I can see the like, I can see a person As if I’m a person Bs doctor, I can see person As private information. That shouldn’t be possible, I think. — C5
@@
The deal here is to treat a certain medical condition and additional data like the first name, last name, birthday, I don’t know, blood type is not required in this case. I basically would only give back treatments or medications that have worked positively for this kind of condition in the past. And the rest is not valuable or not relevant for this situation, I would say. — P5
@@
I mean, I don’t know, the whole function, like the whole function seems to be a bit, not really privacy-protecting. — C8
@@
Technically, it’s possible. And I’m not sure if it’s enough, but at least one has to filter whether for each single user, one has the permissions to read it given by the user, and I would think that the user especially has to allow this kind of use. — C2
4.4 Task 3.2
4.4.1 Source Code Analysis for Task 3.2
Thirteen participants successfully realized the functionality of Task 3.2. Out of those, however, none did so in a privacy-compliant way. Several noteworthy implementation aspects that did not quite fit the scoring scheme will be described below. They are denoted as “Special Case” (SC) in Table III. One participant (P3) misunderstood Task 3.2 and added the email functionality to the doctor’s account. Similarly, P8 did implement this task in a way that did not fit the task description either. Instead of sending it to a list of people, the admin has to type each receiver’s email by hand. Neither E1 nor E9 did finish the task from a functional point of view. However, the aspects implemented by either highlight that both wanted to check user consent prior to sending e-mails. E9 added a flag to the user’s profile page to opt into e-mail notifications, whereas E1 added a filter for user consent prior to sending e-mails.
Overall, for Task 4, five control participants delivered a functioning solution, and three prompted participants implemented a functioning solution and two special cases. Five expert participants delivered a functioning solution for the last task. No participant in the control group or the prompted group delivered a privacy-compliant solution for the fourth task. In contrast, two participants in the expert group were able to implement a compliant solution. While the number of functioning solutions was similar for the groups, only the expert group was able to fulfill the privacy requirements.
4.4.2 In Their Words: Interview Insights for Task 3.2
During Task 3.2, twenty-two participants mentioned consent as the main concept to create a compliant solution:
Yes, the parties have all consented to their email addresses being stored there. The question is whether they have also agreed to allow advertising to be sent out. — E8
@@
That shouldn’t be, that shouldn’t be allowed, I think. Especially that I used email addresses that were for doctor purposes to have for advertisement purposes. — C5
Similar to Task 3.2, five had concerns with the task itself, seeing it as violating the users’ privacy, and were unsure if a compliant solution was possible:
They’ve probably never even, have never even agreed to receive advertisement emails, so I think the entire task is probably faulty. — P6
@@
They are allowed to send the email data. So in this case, I think, this should be okay to send the patients emails. — P9